Driving of casing in the second vent well is complete (visible on Highway 70, northeast of the community). They were able to go down to 199′. If all activities remain on schedule, we may see flaring as early as Thursday or Friday.
Six Geoprobes have been installed on the Triche property and more are currently being installed on Dugas & LeBlanc property. DNR is currently negotiating with land owners to install Geoprobes on residential properties in the immediate area. Pics of the geoprobes were posted on yesterday at http://www.flickr.com/photos/assumptionoep/sets/72157630330363084/. Please note that these pictures are in chronological order and the most recent pics are at the end of the set.
We have benn advised that Texas Brine will resume clean up today. Videos of the sinkhole site were posted on yesterday on http://www.youtube.com/user/assumptionla.
The resident briefing will be held tomorrow at 6:00 p.m., here at the command post.